OpenStudy (dtan5457):

Find all the values of x in the interval 0<=x<360 that satisfy 3cos2x=cosx+2

OpenStudy (dtan5457):

So I have 3cos2x-cosx-2=0 now im not sure

OpenStudy (dtan5457):

@sleepyjess

OpenStudy (sleepyjess):

@Kainui

OpenStudy (dtan5457):

once it's factored it easy..but not sure how to do here. if i substitute cos(x)=u, i still don't see how it could work here

OpenStudy (dtan5457):

this is more than likely gonna turn to a quadratic somehow..

mathslover (mathslover):

Hmm!

mathslover (mathslover):

\(3\cos 2x = \cos x + 2 \) Put \(\cos2x = 2\cos^2x - 1\) \(3(2\cos^2 x - 1) = \cos x + 2\) You get : \(6 \cos^2 x - cos x - 5 = 0 \)

mathslover (mathslover):

Use quadratic equation to solve for cos x.

OpenStudy (dtan5457):

How did you know what to substitute? @mathslover

mathslover (mathslover):

Did you mean by the formula? Well, it is pretty clear that we can solve it by quadratic equation but for that we need degree of the polynomial = 2 And since, cos 2x can be simplified into \(\cos^2 x\) ... with the formula : \[ \cos 2x = 2\cos^2 x - 1\] So, I just used it!

mathslover (mathslover):

Once you start solving such type of questions, you will get used to these techniques! Don't worry.. just keep practicing!